On March 31, 2018, Canseco Plumbing Fixtures purchased equipment for $52,000. Residual value at the end of an estimated four-year service life is expected to be $4,000. The company expects the machine to operate for 12,000 hours.

a. Calculate depreciation expense for 2018 and 2019 using straight line method.
b. Calculate depreciation expense for 2018 and 2019 using sum-of-the-years’-digits method.
c. Calculate depreciation expense for 2018 and 2019 using double-declining balance method.
Answer #1

Assuming year of reporting expenses is from April to March

a. Straight Line Depreciation - Straight line depreciation method assumes depreciation of an asset per year is constant and is computed using the following formula

Straight line depreciation = (Purchase price - Residual value) / Estimated life

= (52000 - 4000) / 4

= 48000 / 4 = 12,000 per year

Depreciation expense for 2018 = $12,000

Depreciation expense for 2019 = $12,000

b. Sum-of-the-years-digits Depreciation

Sum of the years digits Depreciation method is a form of accelerated depreciation wherein the depreciation in the initial years of the life of the asset is greater than the depreciation in the later years.

Estimated life = 4 years

Sum of the year's digits = 1 + 2 + 3 + 4 = 10

Therefore, depreciation in 2018 (year 1) = 4/10 = 0.4 * book value of asset

Depreciation expense in 2018 = 0.4 * 52000 = $20,800

depreciation in 2019 (year 2) = 3/10 = 0.3 * book value of asset

Depreciation expense in 2019 = 0.3 * 52000 = $15,600

c. Double-declining balance Depreciation

Double declining balance Depreciation method is another accelerated depreciation method which doubles the depreciation rate computed in the straight line method but this doubled rate is applied on the carrying value of the asset in the balance sheet (as against the purchase price of the asset)

Depreciation rate as per Straight line method = Depreciation per year as per Straight line / (Purchase price - Residual value)

= 12,000 / (52,000 - 4,000)

= 12,000 / 48,000

= 25%

Depreciation rate in Double declining balance = 25% * 2 = 50%

Therefore Depreciation computation is as follows:

Purchase price $52,000
(-) Depreciation for 2018 at 50% ($26,000)
Balance $26,000
(-) Depreciation for 2019 at 50% on balance ($13,000)
Balance $13,000

Therefore, Depreciation expense for 2018 = $26,000

Depreciation expense for 2019 = $13,000